The lenght of rectangular plot is twice of its breadth... nd thd diaoganal of plot is root 9 m.. find its perimeter. .
Answer by JulietG(1812)

Let Length be L
Let Breadth be W (width)
What we're really measuring is a triangle.
In a right triangle, a^2 (which is L^2) + b^2 (width) = c^2 (diagonal)
We're told that L = 2W
and we're told that the diagonal is the square root of 9
.
Plug in the numbers to a^2 + b^2 = c^2
.
2W^2 + W^2 = sqrt 9
Add the W^2, and root the 9
3W^2 = 3
Divide each side by 3
W^2 = 1
.
If w^2 = 1, then w must be 1 (1*1=1)
If w = 1, then L must be 2 ("twice the breadth")
.
Therefore, the sides of your plot are 1 and 2.
To find the perimeter (remember, it's a fence), take 2 pieces of length and 2 pieces of width.
(2*1) + (2*2) = P
2 + 4 = P
6=P